OpenStudy (anonymous):

8x-3y=1 -2x+3y=11 A.(-1,-3) B.(-1,3) C.(2,5) D.(5,2)

OpenStudy (anonymous):

Hello...you having trouble with this one ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

set it up like this.... 8x - 3y = 1 -2x + 3y = 11 -------------add Can you add them ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what did you get when you added them ?

OpenStudy (anonymous):

6 6 12

OpenStudy (anonymous):

no exactly.....lets do this step by step.... 8x - 2x = 6x -3y + 3y = this is where you messed up this equals 0 because they cancel each other out. You cannot forget about the signs. It should equal : 6x + 0 = 12 6x = 12 Now can you solve for x ?

OpenStudy (anonymous):

ohh yeea

OpenStudy (anonymous):

did you get x yet ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do you know how to get x ?

OpenStudy (anonymous):

gtg

OpenStudy (anonymous):

Even though you are gone, I will explain... 6x = 12 (divide both sides by 6 to get x by itself) 6x/6 = 12/6 x = 2 now sub 2 in for x in either of the original equations to get y. If you are still having trouble when you get back......tag me and I will help :)